# The Atlas Model and SDEs with Boundary Interaction

### v1

@@ 3.1 Convergence to the Generalised Mean-Field Limit
The objective of this section is to study McKean–Vlasov SDEs (1.9) and (1.11). We show that both equations are equivalent and that they admit at least one solution. Throughout this section, we shall assume that $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space carrying a Brownian motion $W$. In addition, for a path $f\in C([0,\infty))$, we set $(f)^{\ast}_{t}=\sup_{0\leq s\leq t}f_{s}$ and $\lvert f\rvert^{\ast}_{t}=\sup_{0\leq s\leq t}\lvert f_{s}\rvert$ for $t\geq 0$.
@@
Furthermore, recall that the signed measure $v$ (identified with its density) appearing in (1.10) and which serves as an initial condition for McKean–Vlasov SDE (1.9) is given in terms of $m$ by $v(x)=m([0,x])$ for $x\geq 0$.
@@
We can directly proceed to the proof of Theorem 1.12.
@@
To construct the solution, we shall employ the Schauder fixed-point theorem. To obtain a suitable subset of $C([0,\infty))$, in which to search for the fixed point, we use [3, Theorem 1.12]. It guarantees the existence of a solution to McKean–Vlasov SDE (1.11) in the case where the initial condition $m$ is a finite measure. Let us begin with some preparation. We equip the space $C([0,\infty))$ with the metric
@@
Next, define the finite measure $m_{\ast}$ on $0,\infty)$ through its cumulative distribution function $F_{\ast}\mathpunct{:}[0,\infty)\to[0,\infty)$, given by
$F_{\ast}(x)=\Bigl(\sup_{0\leq y\leq x}m([0,y])\Bigr)_{+}$
for $x\geq 0$. It holds that $F_{\ast}(0)<\frac{1}{\alpha}$ and $0\leq F_{\ast}(x)\leq\frac{1}{\alpha}$ for $x\geq 0$. If $F_{\ast}(x)=0$ for all $x\geq 0$, we replace $F_{\ast}$ by $x\mapsto\frac{1}{\alpha}\mathbf{1}_{x\geq 1}$. The precise functional form does not matter, we only want to avoid a situation, where $m_{\ast}$ has zero mass. In any case, $m_{\ast}$ is a measure with total mass $\lim_{x\to\infty}F_{\ast}(x)\leq\frac{1}{\alpha}$. Thus, by [3, Theorem 1.12], McKean–Vlasov SDE (1.11) has a nondecreasing continuous unique solution $\ell^{\ast}=(\ell^{\ast}_{t})_{t\geq 0}$ with initial condition $m_{\ast}$. Note that Theorem 1.12 in [3] is phrased in terms of initial conditions with unit mass. To switch back and forth between McKean–Vlasov SDE (1.11) and the equation from [3, Theorem 1.12], we simply replace $m_{\ast}$ and $\ell^{\ast}$ by $m_{\ast}/m_{\ast}([0,\infty))$ and $\ell^{\ast}/m_{\ast}([0,\infty))$, respectively. Then, the coefficient appearing in front of $\ell$ in the equation from [3] becomes $\alpha m_{\ast}([0,\infty))\leq 1$. If the inequality is strict, we are in the subcritical regime, otherwise in the critical one. In either case, [3, Theorem 1.12] indeed guarantees the existence of a unique global-in-time solution.
With this at hand, let us introduce the map $\Lambda\mathpunct{:}C([0,\infty))\to C([0,\infty))$ given by
@@
for $t\geq 0$. Note that by Lemma 4.2, the integral on the right-hand side is finite for any $f\in C([0,\infty))$ and $t\geq 0$. Next, we claim that for any $f\in C([0,\infty))$ with $f_{t}\leq\ell^{\ast}_{t}$ for $t\geq 0$, it holds that $\Lambda_{t}(f)\leq\ell^{\ast}_{t}$. Indeed, applying Lemma 4.2 twice, we see that
@@
$\leq\int_{0}^{\infty}\mathbb{P}\biggl(\inf_{0\leq s\leq t}(x+W_{s}-\alpha\ell^{\ast}_{s})\leq 0\biggr)m_{\ast}([0,x])\,\mathrm{d}x$
@@
$=\int_{[0,\infty)}\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha\ell^{\ast}_{s})_{-}\biggr]\,\mathrm{d}m_{\ast}(x)$
$=\ell^{\ast}_{t}$
for $t\geq 0$, as desired. Next, for $f\in C([0,\infty))$, let $\lVert f\rVert_{\text{TV},[0,t]}$ denote the total variation of $f|_{[0,t]}$. Then, for $t\geq 0$ and any $f\in C([0,\infty))$ with $(f)^{\ast}_{t}\leq\ell^{\ast}_{t}$, it holds that
@@
$\lVert\Lambda(f)\rVert_{\text{TV},[0,t]}$ $\leq\int_{[0,\infty)}\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha f_{s})_{-}\biggr]\,\mathrm{d}\lvert m\rvert(x)$
@@
$\leq\int_{[0,\alpha\ell^{\ast}_{t}]}\bigl(\mathbb{E}[\lvert W_{t}\rvert]+\alpha\ell^{\ast}_{t}\bigr)\,\mathrm{d}\lvert m\rvert(x)+\int_{(\alpha\ell^{\ast}_{t},\infty)}\frac{\sqrt{2t}}{\sqrt{\pi}}e^{\frac{3(\alpha\ell^{\ast}_{t})^{2}}{2t}}e^{-\frac{x^{2}}{4t}}\,\mathrm{d}\lvert m\rvert(x)$
$<\infty,$ (4.5)
where we inserted the estimate from (4.1).
Let us now define $\mathcal{K}$ to be the set of $f\in C([0,\infty))$ such that $f_{t}\leq\ell^{\ast}_{t}$ for $t\geq 0$. Since $\Lambda_{t}(f)\leq\ell^{\ast}_{t}$, $t\geq 0$, for all $f\in\mathcal{K}$, the function $\Lambda$ maps $\mathcal{K}$ into itself. Next, (4.2) shows that $\Lambda(\mathcal{K})$ is contained in a compact subset of $\mathcal{K}$. Clearly, $\mathcal{K}$ is a nonempty convex and closed subset of $C([0,\infty))$. Hence, if $\Lambda$ is continuous on $\mathcal{K}$, it follows from the Schauder fixed-point theorem that $\Lambda$ has a fixed point in $\mathcal{K}$. Showing continuity of $\Lambda$ is straightforward. Fix a sequence $(f^{k})_{k\geq 1}$ in $\mathcal{K}$ which converges to $f\in\mathcal{K}$. Fix $n\geq 1$. Then, for any $k\geq 1$ and $x\geq 0$, it holds that
$\biggl\lvert\mathbb{E}\biggl[\sup_{0\leq t\leq n}(x+W_{t}-\alpha f_{t})_{-}\biggr]-\mathbb{E}\biggl[\sup_{0\leq t\leq n}(x+W_{t}-\alpha f^{k}_{t})_{-}\biggr]\biggr\rvert\leq\alpha\lvert f-f^{k}\rvert^{\ast}_{n}.$
Thus, using a similar decomposition as in (4.2), we find
@@
$\lvert\Lambda(f)-\Lambda(f^{k})\rvert^{\ast}_{n}\leq\alpha\lvert m\rvert([0,z])\lvert f-f^{k}\rvert^{\ast}_{n}+2\int_{(z,\infty)}\frac{\sqrt{2n}}{\sqrt{\pi}}e^{\frac{3(\alpha\ell^{\ast}_{n})^{2}}{2n}}e^{-\frac{x^{2}}{4n}}\,\mathrm{d}\lvert m\rvert(x)$
@@
for $z\geq\alpha\ell^{\ast}_{n}$. Now, given any $\epsilon>0$, we first choose $z$ sufficiently large such that the second term on the right-hand side above is smaller than $\frac{\epsilon}{2}$. Then, we let $K\geq 1$ be large enough such that for all $k\geq K$, it holds that
$\lvert f-f^{k}\rvert^{\ast}_{n}\leq\frac{\epsilon}{2\alpha\lvert m\rvert([0,z])}.$
@@
Together, this implies that $\lvert\Lambda(f)-\Lambda(f^{k})\rvert^{\ast}_{n}\leq\epsilon$ for all $k\geq K$. Since $\epsilon>0$ war arbitrary, we obtain that $\lvert\Lambda(f)-\Lambda(f^{k})\rvert^{\ast}_{n}\to 0$ as $k\to\infty$. This holds for any $n\geq 1$, so it finally follows that $\Lambda(f^{k})\to\Lambda(f)$ as $k\to\infty$ in $C([0,\infty))$. This concludes the proof. ∎

### v2

@@ 3.1 Convergence to the Generalised Mean-Field Limit
The objective of this section is to study McKean–Vlasov SDEs (1.9) and (1.11). We show that both equations are equivalent and that they admit at least one solution. Throughout this section, we shall assume that $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space carrying a Brownian motion $W$. Moreover, we shall use the convention that $[0,T[=[0,T]$ if $T\in[0,\infty)$ and $0,T[\penalty 10000\ =[0,\infty)$ if $T=\infty$. Lastly, for a path $f\in C([0,T[)$, we set $(f)^{\ast}_{t}=\sup_{0\leq s\leq t}f_{s}$ and $\lvert f\rvert^{\ast}_{t}=\sup_{0\leq s\leq t}\lvert f_{s}\rvert$ for $t\in[0,T[$.
@@
Furthermore, recall that the signed measure $v$ (identified with its density) appearing in (1.10), which serves as an initial condition for McKean–Vlasov SDE (1.9), is given in terms of $m$ by $v(x)=m([0,x])$ for $x\geq 0$.
@@
The following lemma shows that Assumption 1.11 is propagated under the dynamics of McKean–Vlasov SDE (1.11).
Lemma 4.3.
Let Assumption 1.11 be satisfied. Let $f\in C([0,\infty))$, denote the solution to the Skorokhod problem for x+Wfx+W-f by $(X^{x},L^{x})$, and define the signed measure $\tilde{m}$ on $0,\infty)$ by
$\tilde{m}(A)=\int_{[0,\infty)}\mathbb{P}(X^{x}_{t}\in A)\,\mathrm{d}m(x)$
for Borel measurable $A\subset[0,\infty)$ and some $t\geq 0$. Then $\tilde{m}$ satisfies Assumption 1.11.
Proof.
Let us prove the properties stated in Assumption 1.11 in sequence. Local finiteness of $\tilde{m}$ is clear, so we start with $\tilde{m}(\{0\})<\frac{1}{\alpha}$. Note that the map $0,\infty)\ni x\mapsto\mathbb{P}(X^{x}_{t}=0)$ is nonincreasing, since if $\varrho_{x}=\inf\{s\in[0,t]\mathpunct{:}X^{x}_{s}=0\}$ occurs before or at time $t$, then $X^{x}_{t}=X^{0}_{t}$, and $X^{x}_{t}>0$ otherwise, so that
$\mathbb{P}(X^{x}_{t}=0)=\mathbb{P}\bigl(X^{x}_{t}=0,\,\varrho_{x}\leq t\bigr)=\mathbb{P}\bigl(X^{0}_{t}=0,\,\varrho_{x}\leq t\bigr).$
Set $P(x)=\mathbb{P}(X^{0}_{t}=0)-\mathbb{P}(X^{x}_{t}=0)=\mathbb{P}(X^{0}_{t}=0,\,\varrho_{x}>t)$ and $\rho(x)=\frac{1}{\alpha}-m([0,x])$ for $x\geq 0$, so integration by parts yields
$\tilde{m}(\{0\})$ $=\int_{[0,\infty)}\mathbb{P}(X^{x}_{t}=0)\,\mathrm{d}m(x)$
$=\int_{[0,\infty)}\mathbb{P}(X^{0}_{t}=0)\,\mathrm{d}m(x)-\int_{[0,\infty)}P(x)\,\mathrm{d}m(x)$
$=\mathbb{P}(X^{0}_{t}=0)m([0,\infty))+\bigl(P(x)\rho(x)\bigr)\big|_{0}^{\infty}-\int_{[0,\infty)}\rho(x)\,\mathrm{d}P(x)$
$\leq\frac{1}{\alpha}\mathbb{P}(X^{0}_{t}=0)$
$<\frac{1}{\alpha},$
where we used in the fourth line that $\lim_{x\to\infty}\rho(x)=\frac{1}{\alpha}-m([0,\infty))$ and that $x\mapsto P(x)$ is nondecreasing while $\rho(x)\geq 0$. The last inequality follows since $\mathbb{P}(X^{0}_{t}=0)<1$.
Let us now deduce the second property, namely $\tilde{m}([0,x])\leq\frac{1}{\alpha}$ for $x\geq 0$. Note that $y\mapsto X^{y}_{t}$ is a.s. continuous and nondecreasing, so for its generalised inverse $Z_{x}=\inf\{x>0\mathpunct{:}X^{y}_{t}>x\}$, we have $X^{y}_{t}\leq x$ if and only if $y\leq Z_{x}$. Consequently, it holds for any $x\geq 0$ that
$\tilde{m}([0,x])=\int_{[0,\infty)}\mathbb{P}\bigl(X^{y}_{t}\in[0,x]\bigr)\,\mathrm{d}m(y)=\mathbb{E}\biggl[\int_{[0,\infty)}\mathbf{1}_{\{y\leq Z_{x}\}}\,\mathrm{d}m(y)\biggr]=\mathbb{E}\bigl[m([0,Z_{x}])\bigr]\leq\frac{1}{\alpha},$
where we applied Fubini’s theorem separately to the positive and negative part of $m$ in the second equality. This gives the third property, so it remains to verify that (1.13) holds for $\tilde{m}$. We have for $c>0$ that
$\int_{[0,\infty)}e^{-cx^{2}}\,\mathrm{d}\lvert\tilde{m}\rvert(x)\leq\int_{[0,\infty)}\mathbb{E}\bigl[\exp(-c\lvert X^{x}_{t}\rvert^{2})\bigr]\,\mathrm{d}\lvert m\rvert(x).$
The finiteness of the integral on the right-hand side is then a consequence of the fact that
$\lvert X^{x}_{t}\rvert^{2}\geq\frac{1}{2}\lvert x\rvert^{2}-\bigl\lvert W_{t}-f_{t}+L^{x}_{t}\bigr\rvert^{2}\geq\frac{1}{2}\lvert x\rvert^{2}-2\lvert W_{t}-f_{t}\rvert^{2}-2\lvert L_{t}\rvert^{2}$
and that (1.13) is true for $m$ by assumption. ∎
We can now proceed to the proof of Theorem 1.13.
@@
The construction of a solution $\ell\in C([0,\infty))$ proceeds in several step. First, we employ the Schauder fixed-point theorem to establish the existence of local-in-time solution $\ell\in C([0,T])$ for some $T>0$. Then, we use Zorn’s lemma to obtain a (possibly local-in-time) solution to McKean–Vlasov SDE (1.11) defined on a maximal time domain $0,T_{\ast})$ for $T_{\ast}\in(0,\infty]$. Finally, arguing by contradiction, we prove that $T_{\ast}=\infty$.
Local existence: As mentioned above, the local-in-time solution is constructed via the Schauder fixed-point theorem. To obtain a suitable subset of $C([0,T])$ (with $T>0$ to be determined below), in which to search for the fixed point, we use [undefb, Theorem 1.12]. It guarantees the existence of a solution to McKean–Vlasov SDE (1.11) in the case where the initial condition is a finite measure.
Let us begin with some preparation. We equip the space $C([0,\infty))$ with the metric
@@
The space $C([0,T])$ for $T>0$ is endowed with the supremum norm and the induced metric. We introduce the map $\Lambda\mathpunct{:}C([0,\infty))\to C([0,\infty))$ given by
@@
for $t\geq 0$. Note that by Lemma 4.2, the integral on the right-hand side is finite for any $f\in C([0,\infty))$ and $t\geq 0$. In what follows, we will also consider the restriction of $\Lambda$ to $C([0,T])$ for $T>0$ and, slightly abusing notation, denote this mapping by $\Lambda$ as well.
[…]
@@
$\leq\int_{0}^{\infty}\mathbb{P}\biggl(\inf_{0\leq s\leq t}(x+W_{s}-\alpha\tilde{\ell}_{s})\leq 0\biggr)\tilde{m}([0,x])\,\mathrm{d}x$
$=\int_{[0,\infty)}\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha\tilde{\ell}_{s})_{-}\biggr]\,\mathrm{d}\tilde{m}(x)$
$=\tilde{\ell}_{t}$ (4.5)
for any $t\geq 0$, as desired. Throughout the proof, we shall frequently make use of this comparison property.
Next, we wish to define a suitable subset of the continuous functions in which to look for a fixed point. This requires further preparation. As hinted at earlier, by [undefb, Theorem 1.12], for any finite measure $\tilde{m}$ on $0,\infty)$, McKean–Vlasov SDE (1.11) has a unique nondecreasing continuous maximal solution $\tilde{\ell}=(\tilde{\ell}_{t})_{t\in[0,\tilde{T}[}$ with initial condition $\tilde{m}$ for $\tilde{T}\in(0,\infty]$. Note that Theorem 1.12 in [undefb] is phrased in terms of initial conditions with unit mass. To switch back and forth between McKean–Vlasov SDE (1.11) and the equation from [undefb, Theorem 1.12], we simply replace $\tilde{m}$ and $\tilde{\ell}$ by $\tilde{m}/\tilde{m}([0,\infty))$ and $\tilde{\ell}/\tilde{m}([0,\infty))$, respectively. Then, the coefficient appearing in front of $\ell$ in the equation from [undefb] becomes $\alpha\tilde{m}([0,\infty))\in[0,\infty)$. If $\alpha\tilde{m}([0,\infty))<1$ or $\alpha\tilde{m}([0,\infty))=1$ and $\alpha\tilde{m}(\{0\})<1$, we are in the subcritical or critical regime, respectively, in which case [undefb, Theorem 1.12] guarantees the existence of a global-in-time solution, i.e. $\tilde{T}$$\infty$. In the supercritical regime $\alpha\tilde{m}([0,\infty))>1$ and $\alpha\tilde{m}(\{0\})<1$, the maximal solution $\tilde{\ell}$ has a finite interval of existence, so $\tilde{T}\in(0,\infty)$.
We shall apply the above result for two distinct choices of the finite initial measure $\tilde{m}$. By the Hahn decomposition theorem, we can find mutually singular measures $m_{+}$ and $m_{-}$ on $0,\infty)$ such that $m_{+}$ or $m_{-}$ have finite mass and $m=m_{+}-m_{-}$. From Assumption 1.11, it follows that $m_{+}$ must have finite mass. Indeed, if $m_{-}$ has finite mass, then since $m([0,x])\leq\frac{1}{\alpha}$ for $x\geq 0$, we find that $m_{+}([0,\infty))\leq\frac{1}{\alpha}+m_{-}([0,\infty))<\infty$, so $m_{+}$ has finite mass too. Next, let us define another finite measure $\bar{m}$ on $0,\infty)$ through its cumulative distribution function $\bar{F}\mathpunct{:}[0,\infty)\to[0,\frac{1}{\alpha}]$, given by
$\bar{F}(x)=\biggl(m_{+}([0,x])+\frac{1}{\alpha}\mathbf{1}_{x\geq 1}\biggr)\land\frac{1}{\alpha}$
for $x\geq 0$. It holds that $\bar{m}(\{0\})=\bar{F}(0)<\frac{1}{\alpha}$, and $0\leq\bar{m}([0,x])=\bar{F}(x)\leq\frac{1}{\alpha}$ and $m([0,x])\leq\bar{F}(x)=\tilde{m}([0,x])$ for $x\geq 0$. By the above, McKean–Vlasov SDE (1.11) has nondecreasing solutions $\ell^{+}=(\ell^{+}_{t})_{t\in[0,T_{+}[}$ and $\bar{\ell}=(\bar{\ell}_{t})_{t\in[0,\infty)}$ for the initial conditions $m_{+}$ and $\bar{m}$, where $T_{+}\in(0,\infty]$. Here we let $T_{+}=\infty$ and $\ell^{+}_{t}=0$ for $t\in[0,\infty)$ if $m_{+}=0$. Note that if $m_{-}$ vanishes, then $m_{+}([0,\infty))\leq\frac{1}{\alpha}$, so $T_{+}=\infty$.
Let us now define $\mathcal{K}$ to be the set of $f\in C([0,T_{+}[)$ such that $f_{t}\leq\bar{\ell}_{t}$ for $t\in[0,T_{+}[$ and $f_{t}-f_{s}\leq\ell^{+}_{t}-\ell^{+}_{s}$ for $s$, $t\in[0,T_{+}[$ with $s\leq t$. Note that we could have also used $\ell^{+}$ as an upper bound for the set $\mathcal{K}$. However, when we later on attempt to obtain a global solution, it is instrumental to have the globally defined upper bound $\bar{\ell}$. Clearly, $\mathcal{K}$ is a nonempty convex and closed subset of $C([0,T_{+}[)$. Next, we will prove that $\Lambda$ maps $\mathcal{K}$ into itself and that the image of $\mathcal{K}$ under $\Lambda$ is a precompact subset of $C([0,T_{+}[)$. Hence, if we further show that $\Lambda$ is continuous on $\mathcal{K}$, it follows from the Schauder fixed-point theorem that $\Lambda$ has a fixed point in $\mathcal{K}$. Let us begin by proving that $\Lambda(\mathcal{K})\subset\mathcal{K}$. Fix $f\in\mathcal{K}$, so that $f_{t}\leq\bar{\ell}_{t}$ for $t\in[0,T_{+}[$ and $f_{t}-f_{s}\leq\ell^{+}_{t}-\ell^{+}_{s}$ for $s$, $t\in[0,T_{+}[$ with $s\leq t$. We have to show that $\Lambda_{t}(f)\leq\bar{\ell}_{t}$ for $t\in[0,T_{+}[$ and $\Lambda_{t}(f)-\Lambda_{s}(f)\leq\ell^{+}_{t}-\ell^{+}_{s}$ for $s$, $t\in[0,T_{+}[$ with $s\leq t$. The former is an immediate consequence of (4.2). Next, by standard comparison results for the Skorokhod problem, it holds for any $x\geq 0$ that
$X^{f,x}_{t}\geq X^{\ell^{+},x}_{t}$
for $t\in[0,T_{+}[$, where $(X^{g,x},L^{g,x})$ is the solution to the Skorokhod problem for $x+W-\alpha g$ with $g\in C([0,T_{+}[)$. Define the measures $m^{\pm}_{t}$ and $\tilde{m}^{\pm}_{t}$, $t\in[0,T_{+}[$, by
$m^{\pm}_{t}(A)=\int_{[0,\infty)}\mathbb{P}\bigl(X^{f,x}_{t}\in A\bigr)\,\mathrm{d}m_{\pm}(x)\quad\text{and}\quad\tilde{m}^{\pm}_{t}(A)=\int_{[0,\infty)}\mathbb{P}\bigl(X^{\ell^{+},x}_{t}\in A\bigr)\,\mathrm{d}m_{\pm}(x)$
for Borel measurable $A\subset[0,\infty)$ and set $m_{t}=m^{+}_{t}-m^{-}_{t}$. Then, for $x\geq 0$, we find $m^{\pm}_{t}([0,x])\leq\tilde{m}^{\pm}_{t}([0,x])$ and $m_{t}([0,x])\leq\tilde{m}^{+}_{t}([0,x])$. Moreover, as in the proof of Lemma 4.3, one can show that $^{\pm}_{t}, m~±\tilde{m}^{\pm}$, and $m_{t}$ satisfy the integrability condition (1.13). Note now that for $s$, $t\in[0,T_{+}[$ with $s\leq t$, we can write
@@
$\Lambda_{t}(f)-\Lambda_{s}(f)$ $=\int_{[0,\infty)}\mathbb{E}\biggl[\sup_{s\leq u\leq t}\Bigl(X^{f,x}_{s}+(W_{u}-W_{s})-\alpha(f_{u}-f_{s})\Bigr)_{-}\biggr]\,\mathrm{d}m(x)$
@@
$=\int_{[0,\infty)}\mathbb{E}\biggl[\sup_{s\leq u\leq t}\Bigl(x+(W_{u}-W_{s})-\alpha(f_{u}-f_{s})\Bigr)_{-}\biggr]\,\mathrm{d}m_{s}(x).$
Similarly, we have that
@@
$\ell^{+}_{t}-\ell^{+}_{s}=\int_{[0,\infty)}\mathbb{E}\biggl[\sup_{s\leq u\leq t}\Bigl(x+(W_{u}-W_{s})-\alpha(\ell^{+}_{u}-\ell^{+}_{s})\Bigr)_{-}\biggr]\,\mathrm{d}\tilde{m}^{+}_{s}(x).$
Thus, proceeding as in (4.2), using that $m_{t}([0,x])\leq\tilde{m}^{+}_{t}([0,x])$ for $x\geq 0$, one can indeed show that $\Lambda_{t}(f)-\Lambda_{s}(f)\leq\ell^{+}_{t}-\ell^{+}_{s}$.
The precompactness of $\Lambda(\mathcal{K})$ can be argued similarly. Another estimate along the lines of (4.2), using that $(m^{+}_{t}+m^{-}_{t})([0,x])\leq(\tilde{m}^{+}_{t}+\tilde{m}^{-}_{t})([0,x])$ for $x\geq 0$, reveals that
@@
$\bigl\lvert\Lambda_{t}(f)-\Lambda_{s}(f)\bigr\rvert\leq\int_{[0,\infty)}\mathbb{E}\biggl[\sup_{s\leq u\leq t}\Bigl(x+(W_{u}-W_{s})-\alpha(\ell^{+}_{u}-\ell^{+}_{s})\Bigr)_{-}\biggr]\,\mathrm{d}\bigl(\tilde{m}^{+}_{s}+\tilde{m}^{-}_{s}\bigr)(x)$
for $s$, $t\in[0,T_{+}[$ with $s\leq t$. The expression on the right-hand side does not depend on $f\in C([0,T_{+}[)$ and converges to zero as $\lvert t-s\rvert\to 0$. Thus, the elements of $\Lambda(\mathcal{K})$ are equicontinuous and have the pointwise upper bound $\bar{\ell}$. This implies that $\Lambda(\mathcal{K})$ is a precompact subset of $C([0,T_{+}[)$.
To apply the Schauder fixed-point theorem, it remains to show that $\Lambda$ is continuous on $\mathcal{K}$. Fix a sequence $(f^{k})_{k\geq 1}$ in $\mathcal{K}$ which converges to $f\in\mathcal{K}$ and let $t\in[0,T_{+}[$. Then, for any $k\geq 1$ and $x\geq 0$, it holds that
$\biggl\lvert\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha f_{s})_{-}\biggr]-\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha f^{k}_{s})_{-}\biggr]\biggr\rvert\leq\alpha\lvert f-f^{k}\rvert^{\ast}_{t}.$
For $x\geq\alpha\bar{\ell}_{t}$, the estimate (4.1) yields the bound
$\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha f_{s})_{-}\biggr]+\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha f^{k}_{s})_{-}\biggr]\leq\frac{2\sqrt{2t}}{\sqrt{\pi}}\exp\biggl(\frac{3(\alpha\bar{\ell}_{t})^{2}}{2t}\biggr)e^{-\frac{x^{2}}{4t}}.$
Thus, we find
@@
$\bigl\lvert\Lambda(f)-\Lambda(f^{k})\bigr\rvert^{\ast}_{t}\leq\alpha\lvert m\rvert([0,z])\lvert f-f^{k}\rvert^{\ast}_{t}+2\int_{(z,\infty)}\frac{\sqrt{2t}}{\sqrt{\pi}}\exp\biggl(\frac{3(\alpha\bar{\ell}_{t})^{2}}{2t}\biggr)e^{-\frac{x^{2}}{4t}}\,\mathrm{d}\lvert m\rvert(x)$
@@
for $z\geq\alpha\bar{\ell}_{t}$. Now, given any $\epsilon>0$, we first choose $z$ sufficiently large such that the second term on the right-hand side above is smaller than $\frac{\epsilon}{2}$. Then, we let $K\geq 1$ be large enough such that for all $k\geq K$, it holds that
$\lvert f-f^{k}\rvert^{\ast}_{t}\leq\frac{\epsilon}{2\alpha\lvert m\rvert([0,z])}.$
@@
Together, this implies that $\lvert\Lambda(f)-\Lambda(f^{k})\rvert^{\ast}_{t}\leq\epsilon$ for all $k\geq K$. Since $\epsilon>0$ war arbitrary, we obtain that $\lvert\Lambda(f)-\Lambda(f^{k})\rvert^{\ast}_{t}\to 0$ as $k\to\infty$. This holds for any $t\in[0,T_{+}[$, so it finally follows that $\Lambda(f^{k})\to\Lambda(f)$ as $k\to\infty$ in $C([0,T_{+}[)$. This concludes the first step.
Existence of maximal solutions: Let $\mathcal{S}$ be the subset of $\cup_{T\in(0,\infty]}(\{T\}\times C([0,T)))$ consisting of couples $(T,\ell)$ such that $\ell=(\ell_{t})_{t\in[0,T)}$ is a (possibly local-in-time) solution to McKean–Vlasov SDE (1.11) with $\ell_{t}\leq\bar{\ell}_{t}$ for $t\in[0,T)$. Note that here we are working with half-open as opposed to closed intervals. By the previous step, the set $\mathcal{S}$ is nonempty. We endow $\mathcal{S}$ with the partial order $\leq$ given by $(T,\ell)\leq(T^{\prime},\ell^{\prime})$ if $T\leq T^{\prime}$ and $\ell_{t}=\ell^{\prime}_{t}$ for $t\in[0,T)$. We claim that every totally ordered subset of $\mathcal{S}$ has an upper bound in $\mathcal{S}$. Let $\mathcal{C}$ be a totally ordered subset of $\mathcal{S}$ and set $T_{\mathcal{C}}=\sup_{(T,\ell)\in\mathcal{C}}T$. Select a sequence $(T_{n},\ell^{n})_{n\geq 1}$ in $\mathcal{C}$ such that $\lim_{n\to\infty}T_{n}=T_{\mathcal{C}}$. We define $\ell^{\mathcal{C}}=(\ell^{\mathcal{C}}_{t})_{t\in[0,T_{\mathcal{C}})}$ by $\ell^{\mathcal{C}}_{t}=\ell^{n}_{t}$ for $t\in[0,T_{\mathcal{C}})$, where $n$ is minimal with $t<T_{n}$. Clearly, $\ell^{\mathcal{C}}$ is a solution to McKean–Vlasov SDE (1.11) on $0,T_{\mathcal{C}})$ and $\ell^{\mathcal{C}}\leq\bar{\ell}_{t}$ for $t\in[0,T_{\mathcal{C}})$, so $(T_{\mathcal{C}},\ell^{\mathcal{C}})$ is an upper bound of $\mathcal{C}$. This allows us to apply Zorn’s lemma, yielding a maximal element $(T_{\ast},\ell^{\ast})$ of $\mathcal{S}$, corresponding to a solution $\ell^{\ast}$ McKean–Vlasov SDE (1.11) that cannot be extended to a contiguous half-open interval strictly including $0,T_{\ast})$.
Global existence: It remains to prove that $T_{\ast}=\infty$. Recall that in the case that $m_{-}$ vanishes, the solution $\ell^{+}$ with initial condition $m=m_{+}$ is already global, so there is nothing to show. Thus, in the following we assume that $m_{-}$ is nontrivial. We argue by contradiction, supposing that $T_{\ast}<\infty$. Since $\ell^{\ast}$ is upper bounded by $\bar{\ell}$, the finiteness of $T_{\ast}$ implies that $\sup_{t\in[0,T_{\ast})}\ell^{\ast}_{t}\leq\sup_{t\in[0,T_{\ast})}\bar{\ell}_{t}=\bar{\ell}_{T_{\ast}}<\infty$. But then it follows from the inequalities
$\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha\ell^{\ast}_{s})_{-}\biggr]\leq\frac{\sqrt{2t}}{\sqrt{\pi}}+\alpha\bar{\ell}_{t}$
for $x\geq 0$ and $t\in[0,T_{\ast})$, and
$\mathbb{E}\biggl[\sup_{0\leq s\leq t}(x+W_{s}-\alpha\ell^{\ast}_{s})_{-}\biggr]\leq\frac{\sqrt{2t}}{\sqrt{\pi}}\exp\biggl(\frac{3(\alpha\bar{\ell}_{t})^{2}}{2t}\biggr)e^{-\frac{x^{2}}{4t}}$
for $x\geq\alpha\bar{\ell}_{t}$ and $t\in[0,T_{\ast})$ that
$\lVert\ell^{\ast}\rVert_{\textup{TV},[0,t]}\leq\lvert m\rvert([0,\alpha\bar{\ell}_{t}])\biggl(\frac{\sqrt{2t}}{\sqrt{\pi}}+\alpha\bar{\ell}_{t}\biggr)+\int_{(\alpha\bar{\ell}_{t},\infty)}\frac{\sqrt{2t}}{\sqrt{\pi}}\exp\biggl(\frac{3(\alpha\bar{\ell}_{t})^{2}}{2t}\biggr)e^{-\frac{x^{2}}{4t}}\,\mathrm{d}\lvert m\rvert(x),$
where $\lVert\cdot\rVert_{\textup{TV},[0,t]}$ denotes the total variation over the interval $[0,t]$. Since the right-hand side is bounded uniformly over $t\in[0,T_{\ast})$, the function $\ell^{\ast}\in C([0,T_{\ast}))$ can be extended to a bounded and continuous function on $0,T_{\ast}]$.
Now, define the signed measure $m_{\ast}$ on $0,\infty)$ by
$m_{\ast}(A)=\int_{[0,\infty)}\mathbb{P}\bigl(X^{\ell^{\ast},x}_{T_{\ast}}\in A\bigr)\,\mathrm{d}m(x)$
[…]